OpenStudy (anonymous):

Instructions: Help please. Given that log 4= 0.6021, log 9= 0.9542 and log 12= 1.0792, evaluate each logarithm. Problem --> log 36

Parth (parthkohli):

\[\log(36) = \log(9 \times 4) = \log9 + \log 4\]

Parth (parthkohli):

You are given what \(\log9\) and \(\log 4\) are . . .

OpenStudy (anonymous):

thank you so much !